# REP: Resource-Efficient Prompting for Rehearsal-Free Continual Learning

### v2

@@ 4.3 Ablation and Additional Study
REP (ours) 75.34$\pm$0.16 3.58$\pm$0.29 2542.43 4.45
L2P 61.57$\pm$0.66 9.73$\pm$0.47 1609.88 14.20
DualPrompt 68.13$\pm$0.49 4.68$\pm$0.20 1507.42 12.30
CODA-Prompt 75.45$\pm$0.56 1.64$\pm$0.10 1853.54 18.83
Competing Methods with Original Setups. For L2P, DualPrompt, and CODA-Prompt, we used the hyperparameters of the original papers, except for the number of iterations per task and batch size. In particular, smaller batch sizes must be adopted to fit into on-device memory constraints. Nonetheless, we present the results of three competing methods on Split ImageNet-R using their original hyperparameters in Table 5. Despite using the ViT-B backbone, they consume significantly more memory than REP due to their large batch sizes.
5 Discussion
Since our study focuses on scenarios wherein a sequence of tasks is presented within learning environments, our method does not determine the curriculum of learning. In general, actively selecting such curricula could enhance the final average accuracy, motivating future work to extend our method for curriculum-based continual learning. Moreover, although using replay buffer may not be feasible in real-world scenarios due to data privacy concerns, on-device CL is already viewed as a privacy-preserving approach, making rehearsal-based REP a promising direction.

### v3

@@ H.2 REP vs Miro
REP-L (ours) 75.34$\pm$0.16 3.58$\pm$0.29 2542.43 4.45
CODA-Prompt-L 74.43$\pm$0.63 3.65$\pm$0.75 6134.51 15.93
OVOR-L 75.40$\pm$0.74 4.17$\pm$0.65 5040.91 7.48
REP-CODA 73.30$\pm$0.93 4.64$\pm$0.89 5459.71 13.71
REP-OVOR 73.99$\pm$0.84 4.94$\pm$0.94 4486.41 7.48
Figure 8: Energy-accuracy trade-offs for Split CIFAR100 and Split ImageNet-R between REP-L and Miro. To provide a spectrum of energy-accuracy trade-offs, we use pre-trained ResNet-18 (RN18), ResNet-34 (RN34), and ResNet-50 (RN50) for Miro.
When operating within the same energy budget, REP-L consistently outperforms Miro variants, achieving 22–33% higher accuracy across datasets. REP-L demonstrates superior cost-effectiveness, especially on Split ImageNet-R compared to Split CIFAR-100. Both REP-L and Miro prove to be memory-efficient to some extent as they fit comfortably within the on-device memory capacity. Although we explore larger ResNet models for Miro, we do not observe the accuracy levels comparable to REP-L for either dataset. This experiment also underscores the importance of specifically optimizing vision transformers for on-device CL scenarios to advance performance boundaries.
Appendix I Impact of Varying Hyperparameter Values
Guided by the hyperparameter tuning strategies used in DualPrompt [38] and CODA-Prompt [31], we fine-tune key hyperparameters for our ALD approach using REP-L. Our focus is primarily on optimizing the threshold parameter ($\alpha$) and the adjustment factor ($\tau$), which are crucial for ALD’s adaptability and efficiency. We conduct a hyperparameter search using cross-validation, for which we designate 20% of our training dataset as a validation set. Figure 10 summarizes the results. We set the AToM’s merging parameter ($n$) to 8.
We also explore the impact of varying prompt-related hyperparameters, as detailed in Figure 10. Note that computational costs increase with larger prompt sizes and longer prompt lengths, and higher costs do not always translate into improved accuracy. Based on Figure 10, we opt for a prompt pool size of 10 and a prompt length of 5 to strike the right balance between computational efficiency and model accuracy.
Figure 9: Result for ALD’s hyperparameter tuning via cross-validation.
Figure 10: Impact of prompt-related hyperparameters of REP. Note that when we increase the pool size and prompt length, there is a notable increase in computation cost.
Appendix J Additional Related Works
J.1 Adapter-based Rehearsal-free Methods
Considering ADAM [44] as a state-of-the-art method for adapter-based CL, we evaluate all its variations proposed in [44] on Split ImageNet-R with 20 tasks and compare their final average accuracies with REP-L in Table 12. ADAM significantly reduces end-to-end training costs using an adapter model. Unlike REP-L, which performs the backward pass to update prompts, ADAM only updates the FC layers without the backward pass. However, training such a model is memory-intensive and often exceeds on-device memory capacity. Also, when implemented with the pre-trained ViT-L model, ADAM’s capability for generalization seems to be somewhat restricted. This observation does not necessarily indicate a fundamental limitation of ADAM but rather underscores the complexities of using it effectively in its current form.
J.2 Advantages of REP over NAS
While NAS (Neural Architecture Search) can be integrated into REP to drop some layers, it usually involves searching the model from scratch, which is time-consuming. In contrast, our method applies adaptive schedules directly to an existing pre-trained model, significantly saving computation time. So, it can naturally adapt to device-internal resource conditions without relying on external server resources.
J.3 Online or Few-shot CL
In on-device CL, the memory capacity for storing data from new tasks may be limited. Thus, one might consider adopting online or few-shot CL methods to save memory costs, as they require fewer new-task samples to be maintained in memory. Offline CL (our setting) typically yields better accuracy with higher resource usage. However, memory allocation for new samples is not a major constraint in offline CL. All prompt-based methods require 155–269MB to store new samples for both Split CIFAR-100 and Split ImageNet-R, accounting for only 3.4–5.9% of REP’s memory usage. Moreover, modern DL frameworks can store data in storage and proactively prefetch upcoming mini-batches for training, virtually eliminating memory concerns for new samples.
J.4 Prompting in Self-Supervised CL
Although our REP framework is designed for conventional supervised CL, we believe it can naturally extend to self-supervised prompt-based CL methods like HiDe-Prompt [34]. By combining REP ’s resource-efficient techniques with hierarchical prompt optimization, this integration could enable on-device CL for large unlabeled datasets. Such a hybrid approach would preserve the performance benefits of prompt tuning while significantly reducing computational and memory overhead, making self-supervised CL more feasible on resource-constrained devices.
Appendix K REP on Diverse Prompting Methods.
In this experiment, we evaluate the performance and applicability of REP in approaches that directly update a single prompt without relying on prompt selection. CODA-Prompt [31] and OVOR [14] stand out as representative methods. We present experimental results using the Split CIFAR-100 and Split ImageNet-R datasets with a ViT-L backbone in  Table 13. REP offers moderately improved resource efficiency when applied to CODA-Prompt and OVOR. During the prompt update, we were able to apply ALD without mod-ifications, whereas AToM ’s main logic required reimplementation to function with CODA-Prompt and OVOR. This does not indicate a fundamental limitation of REP ’s core design but highlights the need to address this implementation challenge in future work. Nonetheless, we note that REP ’s baseline accuracy and resource efficiency (the first row in  Table 13) significantly outperform those observed in CODA-Prompt and OVOR.